OpenStudy (anonymous):

Given: sec theta= -8/3, tan theta is positive. 1. In what quadrant is the theta located?

OpenStudy (anonymous):

sec=1/cos cos is negative in Q2,Q3 tan is positive in Q1,Q3 Therefore Q3

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

Can you help me with a few more like this?

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

okay I want to feel good about myself

OpenStudy (anonymous):

Do you mind if i upload the picture of the 6 problems i need answers to?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

-8/17 8/15 -15/17 -17/18 15/8 starting from 47

OpenStudy (anonymous):

is the last one -15/8 cause i dont have 15/8

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

OH man I thought that the information given was from the problem you gave on this website. I need a minute dude I need to redo all of them!!!!

OpenStudy (anonymous):

The picture i uploaded is different from the first one.

OpenStudy (anonymous):

Oh, and im not a dude.

OpenStudy (matheducatormcg):

\[\csc \left( \theta \right)=-17/15\] means \[\sin \left(\theta\right)=-15/17\] but problem said \[\cos (\theta)\] was positive, which occurs in the 1st and 4th quadrant. Therefore, 15 is your opposite measurement, 17 is your hypotenuse measurement. The adjacent measurement is given by Pythagorean theorem, \[\sqrt{17^{2}-15^{2}}\] Use this measurement and right triangle trig to find other ratios.

OpenStudy (matheducatormcg):

oh, I forgot to mention that \[\sin (\theta)=-15/17\] but \[\cos (\theta)\] is positive, therefore you are in the 4th quadrant.

OpenStudy (anonymous):

which ones are you working on? the picture above?